OpenStudy (anonymous):

Solve the system by using either the substitution method or the elimination-by-addition method, whichever seems more appropriate. x + y = 1300 .82x + .25y = 667

OpenStudy (anonymous):

elimination method x+y = 1300 3.28x + y = 2668 (after multiplying the 2nd equation by 4) 3.28x + y = 2668 - x + y = 1300 = 2.28x = 1368 then solve for x, sub x to find y